OpenStudy (anonymous):

how do i find ordered pairs for x+y=-2 with x=-7,-2,7

OpenStudy (anonymous):

plug in the Xs in the equation and solve it and for each y with the x makes a coordinate

OpenStudy (whpalmer4):

for example, if x=3 was one of the ones to do (it isn't), you would do x+y=-2 3+y=-2 y=-2-3=-5 (3,-5)